OpenStudy (anonymous):

A box has 10 marbles in it, 2 red and 8 white. Suppose we draw a marble from the box, replace it, and then draw another. Find the probability that (a) Both marbles are red (b) Just one of the two marbles is red

OpenStudy (anonymous):

a) (1/5)(1/5)=1/25 b)(1/5)(4/5)=4/25

OpenStudy (zarkon):

hmmm...not real happy with your answer to (b) ;)

OpenStudy (anonymous):

what do you think it should be?

OpenStudy (zarkon):

8/25

OpenStudy (anonymous):

how did you get that?

OpenStudy (zarkon):

one red that mean either the first is red 2nd is white or the 2nd is red and 1st iw white \[\frac{2}{10}\frac{8}{10}+\frac{8}{10}\frac{2}{10}=\frac{8}{25}\]

OpenStudy (anonymous):

makes sense n since i wasnt sure go with that

OpenStudy (zarkon):

if you find the prob both are white you get 64/100 1/25+8/25+64/100=1 Law of total probability is satisfied